OpenStudy (anonymous):

find the definite integral y = x(1-x) from -2 to 2

OpenStudy (anonymous):

you can multiply the x into the parenthese and obtain:, then its just routine

OpenStudy (anonymous):

integrate and evaluate at the limits

OpenStudy (anonymous):

could you please further elaborate thankx

myininaya (myininaya):

\[\int\limits_{}^{}(x-x^2) dx=\frac{x^{1+1}}{1+1}-\frac{x^{2+1}}{2+1}+C\] \[\int\limits_{a}^{b}f(x) dx=F(x)|_a^b=F(b)-F(a)\]

OpenStudy (anonymous):

man, i caught up with something else, or else i would have explained it to you. But the one above me, did a fairly good job

OpenStudy (anonymous):

thank you @ lagrangeson678 ^.^

OpenStudy (anonymous):

myininaya what did you get as your answer?

myininaya (myininaya):

i found the antiderivative now all you have to do is: (just plug in upper limit) then write minus (then plug in lower limit)

OpenStudy (anonymous):

oh because i got a negative answer

OpenStudy (anonymous):

and according to the solution booklet it is a positive answer i.e 17/3 square units

OpenStudy (anonymous):